A rectangular field of 150 m by 100 m has 10 m a path around it. & Find the area of the path.
1
Expert's answer
2011-05-24T09:24:12-0400
The dimentions of the outer border are A = 150 + 2*10 = 170 m and B = 100 + 2*10 = 120 m. The area of the field with borders is S2 = A*B = 170 * 120 = 20400 m2. The area of the field without the borders is S1 = 150 *100 = 15000m2.
Therefore, the area of the path is
S2 - S1 = 20400 - 15000 = 5400 m2.